Directions: find a common solution for each system of equations. The equations are y=-3x and x+y=6. The answer i got to these equations was (-3,9). Am i correct? If not show how you got your answer. Thanx.
Answer by checkley71(8403)

YOU ARE CORRECT (-3,9)
Y=-3X NOW REPLACE Y BY -3X IN THE OTHER EQUATION.
X-3X=6
-2X=6
X=6/-2
X=-3 ANSWER.
Y=-3*-3
Y=9 ANSWER.